OpenStudy (anonymous):

complete each ordered pair so that it is a equation of the given linear equation x-5y=6 (_,-3) (-4,_)

OpenStudy (whpalmer4):

Numbers in ordered pairs are (x,y) For each pair, plug in the known value, and solve for the unknown. For example, if one of the ordered pairs happened to be (_,5): \[x-5(5) = 6\]\[x-25=6\]\[x-25+25 = 6+25\]\[x=31\]So the answer would be \((31,5)\)